OpenStudy (anonymous):

Select the set of equations that represents the following situation: Mike invested $199. for one year. he invested part of it at 5% and the rest at 7%. At the end of the year he earned $10.75 in interest. How much did Mike invest at each rate of interest? can you please help me

OpenStudy (anonymous):

the list of possible choices are a) x+y=199; 0.05x+0.07y = 10.75 b) 0.05x+0.07y = 199; x+y = 10.75 c) 0.05*199+0.07*199 = 10.75 d) x+y= 199; 5x + 7y = 10.75

OpenStudy (anonymous):

a).

OpenStudy (anonymous):

Thank you so much!